write the rational number as the ratio of two integers in the form A/B and then as a repeating decimal:
-7 1/3= -22/3

Found 2 solutions by stanbon, masheila_b:
Answer by stanbon(48546)

write the rational number as the ratio of two integers in the form A/B and then as a repeating decimal:
-7 1/3=
---------
= -22/3

= -7.3333333333....
Cheers,
Stan H.

Answer by masheila_b(1)

write the rational number as the ratio of two integers in the form A/B and then as a repeating decimal:
-7 1/3= -22/3